# Personalized Enhanced Federated Multi-View Clustering via Heat-Kernel Tensor Decomposition

### v2

@@ 3.5.3 Tensor Distance
We know that, the Euclidean distance between data points $x_{i}$ and cluster centers $a_{k}$ can be represented as:
@@
where $\rho_{lm}$ is the metric coefficient defined by
@@
and $\|q_{1}-q_{m}\|_{2}$ is defined as the location distance between $\mathcal{X}_{g_{1}g_{2}\ldots g_{N}}$ and $\mathcal{A}_{g_{1}g_{2}\ldots g_{N}}$
@@
Here, $\theta$ is a user-defined parameter with $\theta>0$ and $P$ is the metric matrix.
@@ 4 Proposed Methodology
This section presents the proposed heat-kernel enhanced tensorized framework for personalized federated multi-view clustering. The initial step involves formulating the problem setting and introducing the federated kernel Euclidean distance metric. In the following section, the calculation of federated heat-kernel coefficients adapted to the multi-view tensor data structure will be delineated. Finally, the overall algorithmic framework must be described, including local updates, global aggregation, and privacy-preserving mechanisms.
@@ 4.3 Federated Heat-Kernel Coefficients
5 The Enhanced Federated Kernel Multi-View Clustering
In this section, the proposed Enhanced Federated Kernel Multi-View Clustering (E-FKMVC) algorithm will be presented. The E-FKMVC algorithm is a refinement of the traditional multi-view clustering framework that operates within a federated learning environment. This refinement enables multiple clients to collaboratively cluster their local multi-view datasets while preserving data privacy.
@@ 5.1 Personalized Federated Learning
$J_{E-FKMVC}^{(\ell)}=J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})=\sum_{\ell=1}^{M}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{~FKED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})$ (22)
@@
where the updates are performed iteratively until convergence, subject to the normalization constraints in Eqs. 23 and 24.
@@
We establish the necessary optimality conditions for the E-FKMVC objective function using Lagrangian optimization and derive the closed-form update rules for each parameter set. The proof proceeds by applying the method of Lagrange multipliers to the constrained optimization problem and solving the resulting system of equations.
@@
To derive the update equation for membership coefficients $\mu_{[\ell]ik}^{*}$ for client $\ell$, we begin by computing the partial derivative of the Lagrangian in Eq. 25 with respect to $\mu_{[\ell]ik}^{*}$.
Taking the partial derivative with respect to $\mu_{[\ell]ik}^{*}$:
@@
Setting this derivative equal to zero for optimality:
@@
Solving for $\mu_{[\ell]ik}^{*}$:
@@
To determine the Lagrange multiplier $\lambda_{1i}$, we apply the normalization constraint $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$:
@@
Solving for the Lagrange multiplier:
@@
Substituting this back into Eq. 33:
@@
This establishes the membership matrix update rule in Eq. 26.
@@
For cluster center updates $a_{[\ell]kj}^{h}$, we differentiate the objective function in Eq. 22 with respect to $a_{[\ell]kj}^{h}$:
@@
To compute $\frac{\partial\text{FKED}}{\partial a_{[\ell]kj}^{h}}$, we recall that:
@@
Let $\phi_{[\ell]ik}^{h}=\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}$. Then:
@@
$=-2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$ (43)
Substituting this into Eq. 38 and setting equal to zero for optimality:
$v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\left[-2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)\right]=0$ (44)
Simplifying and rearranging:
$\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\delta_{[\ell]ij}^{h}$ $\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})=0$ (45)
Let $w_{[\ell]ijk}^{h}=(\mu_{[\ell]ik}^{*})^{m}v_{[\ell]h}^{\alpha}\exp\left(-\sum_{j^{\prime}=1}^{d_{[\ell]}^{h}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$. Then:
@@
This yields the cluster center update rule in Eq. 27.
@@
For view weight updates $v_{[\ell]h}$, we differentiate the Lagrangian with respect to $v_{[\ell]h}$:
@@
Setting this derivative equal to zero for optimality:
@@
Solving for $v_{[\ell]h}$:
@@
Applying the normalization constraint $\sum_{h=1}^{s(\ell)}v_{[\ell]h}=1$:
@@
Solving for the Lagrange multiplier:
@@
Substituting back:
@@
This establishes the view weight update rule in Eq. 28.
@@
These conditions are satisfied by construction, as the FKED function is continuous and bounded, the probability simplex constraints define compact convex sets, and the update rules yield unique solutions when the denominators are non-zero (guaranteed by the strict positivity of FKED values and the non-degeneracy assumption that no data points are identical to cluster centers).
@@
Algorithm 1 Enhanced Federated Kernel Multi-View Clustering (E-FKMVC)
0: Multi-view data $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ for clients $\ell=1,\ldots,M$, number of clusters $c(\ell)$, fuzziness parameter $m>1$, view weight exponent $\alpha>1$, heat-kernel parameters $\delta_{[\ell]ij}^{h}$
0: Membership matrices $U_{[\ell]}^{*}$, cluster centers $A_{[\ell]}^{h}$, view weights $V_{[\ell]}$
1:Initialization: Randomly initialize $U_{[\ell]}^{*}$, $A_{[\ell]}^{h}$, and $V_{[\ell]}$ for all clients $\ell$
2:repeat
3:  for each client $\ell=1$ to $M$ do
4:   Update membership matrix $U_{[\ell]}^{*}$ using Eq. 26
5:   Update cluster centers $A_{[\ell]}^{h}$ using Eq. 27
6:   Update view weights $V_{[\ell]}$ using Eq. 28
7:  end for
8:until convergence criteria met (e.g., changes in $U_{[\ell]}^{*}$, $A_{[\ell]}^{h}$, and $V_{[\ell]}$ below a threshold)

### v3

@@ 2.2.3 Tensor Distance
In this subsection, we introduce the concept of tensor distance, which extends the traditional Euclidean distance to higher-order tensors. This is particularly relevant in the context of multi-view data represented as tensors, where capturing the relationships between different modes is crucial.
The traditional Euclidean distance between two vectors $x_{i}$ and $a_{k}$ in $\mathbb{R}^{d}$ is defined as
@@
where $\|\cdot\|$ denotes the Euclidean norm. However, when dealing with multi-view data represented as tensors, it is essential to consider the interactions across different modes. To address this, we propose the following definition of tensor distance:
@@
In Eq. 19, the tensor distance $d_{TD}(\mathcal{X},\mathcal{A})$ is computed using a weighted Euclidean distance formulation, formulated in terms of the vectorized representations $x$ and $a$ of the tensors $\mathcal{X}$ and $\mathcal{A}$, respectively. The matrix $P$ is a metric matrix that incorporates the heat-kernel coefficients $\rho_{lm}$, which modulate the contribution of each pair of elements $(x_{l},a_{l})$ and $(x_{m},a_{m})$ to the overall distance. The metric coefficients $\rho_{lm}$ are defined as follows:
@@
In Eq. 20, the term $\|q_{1}-q_{m}\|_{2}$ represents the Euclidean distance between two points $q_{1}$ and $q_{m}$ in the feature space, which can be computed as:
@@
In Eq. 21, $g_{n}$ and $g_{n}^{\prime}$ denote the indices along the $n$-th mode of tensors $\mathcal{X}$ and $\mathcal{A}$, respectively. The parameter $\theta$ in Eq. 20 is a scaling factor that controls the spread of the heat kernel, influencing how rapidly the weights decay with increasing distance between points in the feature space.
@@ 3 Proposed Methodology
This section presents the proposed federated heat-kernel enhanced multi-view clustering framework, which does not involve the sharing of raw data across clients. The section commences with the formulation of the problem and the introduction of the federated kernel Euclidean distance (FKED) metric. Subsequently, the federated heat-kernel coefficients (FedH-KC) are defined, which adaptively weight feature distances in the federated setting. In summary, the overarching objective function for federated multi-view clustering using FKED and FedH-KC is hereby presented.
@@ 3.4 Personalized Federated Learning
$J_{E-FKMVC}^{(\ell)}=J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})=\sum_{\ell=1}^{M}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{~FKED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})$
@@
where the updates are performed iteratively until convergence, subject to the normalization constraints in Eqs. 26 and 27.
@@
We establish the necessary optimality conditions for the E-FKMVC objective function using Lagrangian optimization and derive the closed-form update rules for each parameter set. The proof proceeds by applying the method of Lagrange multipliers to the constrained optimization problem and solving the resulting system of equations. This classical optimization technique allows us to find optimal solutions while respecting the normalization constraints on memberships $U_{[\ell]}^{*}$ and view weights $V_{[\ell]}$.
@@
The membership matrix $U_{[\ell]}^{*}$ determines how strongly each data point $i$ belongs to each cluster $k$. To derive the optimal update equation for membership coefficients $\mu_{[\ell]ik}^{*}$ for client $\ell$, we employ the method of Lagrange multipliers. This approach transforms our constrained optimization problem into an unconstrained one by incorporating the constraint $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$ directly into the objective function.
We begin by computing the partial derivative of the Lagrangian in Eq. 28 with respect to $\mu_{[\ell]ik}^{*}$. This derivative tells us how the objective function changes when we slightly adjust the membership of data point $i$ in cluster $k$:
@@
Here, the power rule of differentiation yields the factor $m(\mu_{[\ell]ik}^{*})^{m-1}$, while $\lambda_{1i}$ is the Lagrange multiplier that enforces the membership normalization constraint for data point $i$.
At the optimal solution, this derivative must equal zero (a necessary condition for optimality). Setting the derivative to zero gives us:
@@
Rearranging this equation to isolate $\mu_{[\ell]ik}^{*}$, we first solve for $(\mu_{[\ell]ik}^{*})^{m-1}$ and then take the (m1)(m-1)-th root:
@@
At this point, $\lambda_{1i}$ remains unknown. To determine its value, we apply the normalization constraint which requires that all membership values for data point $i$ across all clusters must sum to unity (i.e., $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$). Summing Eq. 36 over all clusters $k$:
@@
This equation allows us to solve for the term containing the Lagrange multiplier:
@@
Substituting this expression back into Eq. 36 eliminates the Lagrange multiplier and yields the final closed-form update rule:
@@
This formula has an intuitive interpretation: the membership of data point $i$ in cluster $k$ is inversely proportional to the weighted distance (across all views) between the data point and the cluster center. Data points closer to a cluster center receive higher membership values for that cluster. The negative exponent 1/(m1)-1/(m-1) ensures this inverse relationship, while the denominator normalizes the memberships to sum to one. This establishes the membership matrix update rule in Eq. 29.
@@
The cluster centers $a_{[\ell]kj}^{h}$ represent the prototype or centroid of each cluster in each view. To find optimal cluster centers, we differentiate the objective function in Eq. 25 with respect to $a_{[\ell]kj}^{h}$, which represents the $j$-th feature coordinate of the $k$-th cluster center in view $h$ held by client $\ell$:
@@
The key challenge here is computing the derivative of the FKED function, which involves an exponential of a sum of squared differences. Recall the definition of FKED from Eq. 22:
@@
To simplify the derivation, we introduce a shorthand notation for the argument of the exponential:
$\phi_{[\ell]ik}^{h}=\sum_{j^{\prime}=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}$
This $\phi_{[\ell]ik}^{h}$ represents the heat-kernel weighted squared distance between data point $i$ and cluster center $k$ in view $h$ held by client $\ell$. Using the chain rule of differentiation (which states that the derivative of a composite function equals the derivative of the outer function times the derivative of the inner function):
@@
$=-2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\exp\left(-\sum_{j^{\prime}=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$ (46)
The factor of 2 arises from differentiating the squared term, and the negative sign comes from differentiating $(x-a)^{2}$ with respect to $a$.
Substituting this result into Eq. 41 and setting the derivative equal to zero (the optimality condition):
$v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\left[-2\delta_{[\ell]ij}^{h}(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})\exp\left(-\sum_{j^{\prime}=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)\right]=0$ (47)
Since $v_{[\ell]h}^{\alpha}>0$, $\delta_{[\ell]ij}^{h}>0$, and the exponential term is always positive, we can simplify by dividing out these non-zero factors:
$\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}$ $\exp\left(-\sum_{j^{\prime}=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)(x_{[\ell]ij}^{h}-a_{[\ell]kj}^{h})=0$ (48)
To make the subsequent algebra more transparent, we define a weight term that combines membership, view importance, and heat-kernel influence:
$w_{[\ell]ijk}^{h}=(\mu_{[\ell]ik}^{*})^{m}v_{[\ell]h}^{\alpha}\exp\left(-\sum_{j^{\prime}=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij^{\prime}}^{h}(x_{[\ell]ij^{\prime}}^{h}-a_{[\ell]kj^{\prime}}^{h})^{2}\right)$
This weight $w_{[\ell]ijk}^{h}$ captures how much influence data point $i$ should have on determining cluster center $k$ in view $h$ for client $\ell$. Points with higher membership values and those closer to the current cluster center receive higher weights. Using this notation:
@@
This result is a weighted average of the data points, where the weights depend on both the membership values and the current distances to the cluster center. This form is characteristic of expectation-maximization style algorithms and ensures that the cluster center moves toward regions of high membership density. This yields the cluster center update rule in Eq. 30.
@@
The view weights $v_{[\ell]h}$ control the relative importance of each view in the clustering process. Views that produce more coherent clustering results should receive higher weights. To derive the optimal view weight update, we differentiate the Lagrangian with respect to $v_{[\ell]h}$:
@@
Note that only the term corresponding to view $h$ survives after differentiation, since the other terms do not depend on $v_{[\ell]h}$. The power rule yields the factor $\alpha v_{[\ell]h}^{\alpha-1}$.
Setting this derivative equal to zero at the optimum:
@@
The derivation now follows the same pattern as the membership update. Solving for $v_{[\ell]h}$:
@@
To determine $\lambda_{2}$, we apply the normalization constraint $\sum_{h=1}^{s(\ell)}v_{[\ell]h}=1$, which ensures that view weights form a valid probability distribution:
@@
Solving for the Lagrange multiplier term:
@@
Substituting back to eliminate the Lagrange multiplier:
@@
This formula assigns higher weights to views with smaller total weighted distances (i.e., views where data points are closer to their assigned cluster centers). The negative exponent $1/(\alpha-1)$ creates this inverse relationship, ensuring that more informative views—those producing tighter, more coherent clusters—receive higher importance weights. This establishes the view weight update rule in Eq. 31.
@@
These conditions are satisfied by construction, as the FKED function is continuous and bounded, the probability simplex constraints define compact convex sets, and the update rules yield unique solutions when the denominators are non-zero (guaranteed by the strict positivity of FKED values and the non-degeneracy assumption that no data points are identical to cluster centers). ∎
3.4.1 Aggregation of Local Models
After each client $\ell$ updates its local parameters $U_{[\ell]}^{*}$, $A_{[\ell]}^{h}$, and $V_{[\ell]}$ using the derived update rules, these local models can be aggregated at a central server to form a global clustering model. The aggregation process involves averaging the cluster centers and view weights across all clients, weighted by the number of samples at each client. The global cluster centers and view weights are computed as follows:
$A_{global}^{h}=\frac{1}{N}\sum_{\ell=1}^{M}n(\ell)A_{[\ell]}^{h}$ (60)
$V_{global}=\frac{1}{N}\sum_{\ell=1}^{M}n(\ell)V_{[\ell]}$ (61)
where $N=\sum_{\ell=1}^{M}n(\ell)$ is the total number of samples across all clients. The global membership matrix can be constructed by concatenating the local membership matrices from each client.
3.4.2 Personalization of Global Model
Once the global model parameters $A_{global}^{h}$ and $V_{global}$ are obtained, they can be sent back to each client for personalization. Each client $\ell$ can update its local cluster centers and view weights by incorporating the global parameters, allowing for a personalized clustering model that leverages both local data characteristics and global insights. The personalization step involves setting:
$A_{[\ell]}^{h}=A_{global}^{h}-\epsilon\cdot\nabla_{A_{[\ell]}^{h}}J_{E-FKMVC}^{(\ell)}$ (62)
$V_{[\ell]}=V_{global}-\epsilon\cdot\nabla_{V_{[\ell]}}J_{E-FKMVC}^{(\ell)}$ (63)
where $\epsilon$ is a small learning rate that controls the extent of personalization based on local data and $\nabla_{A_{[\ell]}^{h}}J_{E-FKMVC}^{(\ell)}$ and $\nabla_{V_{[\ell]}}J_{E-FKMVC}^{(\ell)}$ represent the gradients of the local objective function with respect to the cluster centers and view weights, respectively. To avoid negative weights, the updated view weights are normalized to ensure they sum to one or set to a minimum threshold.
3.4.3 E-FKMVC Algorithm Summary
@@
Algorithm 1 Efficient Federated Kernel Multi-View Clustering (E-FKMVC)
0: Multi-view data $\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ for clients $\ell=1,\ldots,M$, clusters $c(\ell)$, fuzzifier $m>1$, view exponent $\alpha>1$, convergence threshold $\epsilon>0$, max rounds $T$, max local iterations $E$
0: Membership matrices $\{U_{[\ell]}^{*}\}_{\ell=1}^{M}$, cluster centers $\{A_{[\ell]}^{h}\}_{\ell=1}^{M}$, view weights $\{V_{[\ell]}\}_{\ell=1}^{M}$
1:/* Phase 1: Server Initialization */
2: Initialize global cluster centers $A_{global}^{h}$ and view weights $V_{global}$ randomly
3: Broadcast $A_{global}^{h}$, $V_{global}$ to all clients
4:for round $t=1$ to $T$ do
5:  /* Phase 2: Local Client Computation (Parallel) */
6:  for each client $\ell=1$ to $M$ in parallel do
7:   Step 2.1: Heat-Kernel Coefficient Computation
8:   for $h=1$ to $s(\ell)$, $j=1$ to $d_{h}^{(\ell)}$ do
9:    Compute $\delta_{[\ell]ij}^{h}$ using Eq. 23 or Eq. 24
10:   end for
11:   Step 2.2: Local Optimization
12:   for epoch $e=1$ to $E$ do
13:    Update $\mu_{[\ell]ik}^{*}$ $\forall i,k$ using Eq. 29
14:    Update $a_{[\ell]kj}^{h}$ $\forall k,j,h$ using Eq. 30
15:    Update $v_{[\ell]h}$ $\forall h$ using Eq. 31
16:    if local convergence achieved then
17:     break
18:    end if
19:   end for
20:   Step 2.3: Compute and Send Statistics
21:   Compute local objective $J_{E-FKMVC}^{(\ell)}$ using Eq. 25
22:   Send $\{A_{[\ell]}^{h},V_{[\ell]},J_{E-FKMVC}^{(\ell)},n(\ell)\}$ to server
23:  end for
24:  /* Phase 3: Server Aggregation */
25:  Compute client weights: $\omega_{\ell}\leftarrow n(\ell)/\sum_{\ell^{\prime}=1}^{M}n(\ell^{\prime})$
26:  Aggregate global centers: $A_{global}^{h}\leftarrow\frac{1}{N}\sum_{\ell=1}^{M}n(\ell)A_{[\ell]}^{h}$ {Eq. 60}
27:  Aggregate global view weights: $V_{global}\leftarrow\frac{1}{N}\sum_{\ell=1}^{M}n(\ell)V_{[\ell]}$ {Eq. 61}
28:  Broadcast $A_{global}^{h}$, $V_{global}$ to all clients
29:  /* Phase 4: Personalization */
30:  for each client $\ell=1$ to $M$ in parallel do
31:   $A_{[\ell]}^{h}\leftarrow A_{global}^{h}-\epsilon_{A}\cdot\nabla_{A_{[\ell]}^{h}}J_{E-FKMVC}^{(\ell)}$ {Eq. 62}
32:   $V_{[\ell]}\leftarrow\text{Normalize}\left(V_{global}-\epsilon_{V}\cdot\nabla_{V_{[\ell]}}J_{E-FKMVC}^{(\ell)}\right)$ {Eq. 63}
33:  end for
34:  /* Phase 5: Global Convergence Check */
35:  $J_{global}^{(t)}\leftarrow\sum_{\ell=1}^{M}\omega_{\ell}J_{E-FKMVC}^{(\ell)}$
36:  if $|J_{global}^{(t)}-J_{global}^{(t-1)}|/|J_{global}^{(t-1)}|<\epsilon$ then
37:   break {Global convergence achieved}
38:  end if
39:end for
40:return $\{U_{[\ell]}^{*},A_{[\ell]}^{h},V_{[\ell]}\}_{\ell=1}^{M}$